OpenStudy (anonymous):

the function f described by f(x)=2x^2+3x-2 find the function of f(-9)=

jhonyy9 (jhonyy9):

f(-9)=2(81)+3*9-2=162+27-2=162+25=187

OpenStudy (anonymous):